If the probability of an event is 0.15, what are the odds against the event? Write answer as a ratio using a colon to separate the two numbers, and the ratio must be reduced to the lowest terms.
Answer by Theo(13342)

if the probability for an event is .15, then the probability against the event is (1 - .15) = .85.
the odds for the event is the probability for divided by the probability against = .15 / .85 = 15 / 85 = 3 / 17 = 3:17.
the odds against the event is the probability against divided by the probability for = .85 / .15 = 85 / 15 = 17 / 3 = 17:3.

here's a reference.

https://stats.seandolinar.com/statistics-probability-vs-odds/